OpenStudy (anonymous):

PLEZZZZ SHOW ME 3/4 M =5/6 PLEZZ SHOW WORK

OpenStudy (anonymous):

Divide both sides by 3/4

OpenStudy (anonymous):

20/18=5/3

OpenStudy (anonymous):

no sorryM= 10/9

OpenStudy (anonymous):

3/4m = 5/6 3/4m divided by 3/4 = 5/6 divided by 3/4 m = 5/6 * 4/3 m = 20/18 m = 10/9 m = 1 1/9

OpenStudy (anonymous):

*m = 1 and 1/9

OpenStudy (mertsj):